.Whichofthefolowingistherectangularform ofofthecomplexnumber7.2e

a)0.509+i7.182 b)0.509-2.281c)2.182+i7.217 d)0.217+i7.7.182


1
Expert's answer
2021-03-11T06:19:29-0500

a) "z=0.509+i7.182"

"|z|=\sqrt{(0.509)^2+(7.182)^2}\approx7.200"

"\theta=\tan^{-1}(\dfrac{7.182}{0.509})\approx1.500"


"z=7.200e^{i1.500}"

b) "z=0.509-i2.281"

"|z|=\sqrt{(0.509)^2+(-2.281)^2}\approx2.337"

"\theta=\tan^{-1}(\dfrac{-2.281}{0.509})\approx-1.351"


"z=2.337e^{-i1.351}\ or\ z=2.337e^{i4.932}"


c) "z=2.182+i7.217"

"|z|=\sqrt{(2.182)^2+(7.217)^2}\approx7.540"

"\theta=\tan^{-1}(\dfrac{7.217}{2.182})\approx1.277"


"z=7.540e^{i1.277}"

d) "z=0.217+i7.182"

"|z|=\sqrt{(0.217)^2+(7.182)^2}\approx7.185"

"\theta=\tan^{-1}(\dfrac{7.182}{0.217})\approx1.541"


"z=7.185e^{i0.1541}"

Answer

a) "z=0.509+i7.182=7.200e^{i1.500}"